# IntroStyle: Training-Free Introspective Style Attribution using Diffusion Features

### v1

@@ Appendix F Results on WikiArt and SHacks Datasets
Figure 17 and Figure 18 show retrieval results of the proposed IntroStyle method on WikiArt. Remarkably, IntroStyle is able to identify images of the same style despite a great diversity of content. Note that the images retrieved rarely present images of similar content, indicating the robustness of the proposed approach in focusing on styles. Furthermore, we show more retrieval results in Figure 19 for the SHacks Dataset, where our method works effectively and achieves high retrieval accuracy as well.
The retrieval results, in fact, also demonstrate the effectiveness of the hacked prompts, where images generated by the hacked prompts are indeed similar to the reference image in styles. Our approach successfully retrieves images generated by the hacked prompts, which makes it potentially a strong method for protecting artist styles in real-world applications.
c.1 Katsushika Hokusai’s The Great Wave off Kanagawa
c.2 Design an ukiyo-e style image of a colossal ocean wave with finger-like tendrils of foam. Depict fishermen in traditional boats beneath the wave, and a serene mountain peak on the horizon.
c.3 Design an image of a colossal ocean wave with foam.
c.1 Vincent Van Gogh’s Starry Night
c.2 Design a post-impressionist nightscape with swirling stars and a crescent moon over a small village.
c.3 A small village with crescent moon and stars at night sky
Figure 16: Style Hacks (SHacks) dataset samples. Each row shows images generated for a reference image by an artist. From left to right: reference image, three images generated by “hacked” prompts, and three images generated by “innocent” prompts.
@@
Figure 19: Additional Image Retrieval on SHacks Dataset for IntroStyle with images ranked highest to lowest from left to right. Green colors indicate correct and red for incorrect retrievals.

### v2

@@ Appendix E Results on WikiArt and ArtSplit Datasets
Figs. 13 and 14 show retrieval results of the proposed IntroStyle method on WikiArt and Fig 12 compares our retrievals with CSD. Remarkably, IntroStyle can identify images of the same style despite a great diversity of content. Note that the images retrieved rarely present images of similar content, indicating the robustness of the proposed approach in focusing on styles. Furthermore, we show more retrieval results in Figs. 15 and 16 for the ArtSplit Dataset, where our method works effectively and achieves high retrieval accuracy over state-of-the-art method (CSD).
(a) Queries
(b) IntroStyle (ours)
(c) CSD [38]
Figure 12: Image Retrieval Results on WikiArt Dataset for IntroStyle, with images ranked highest to lowest from left to right compared with CSD. Green colors indicate correct and red for incorrect retrievals.
@@
(c) CSD [38]
Figure 15: Additional Image Retrieval Results for Style-based Evaluation on ArtSplit Dataset, with images ranked highest to lowest from left to right. The images are filtered to a fixed semantic. Green colors indicate correct and red for incorrect retrievals.
(a) Queries
(b) IntroStyle (ours)
(c) CSD [38]
Figure 16: Additional Image Retrieval Results for Semantic-based Evaluation on ArtSplit Dataset, with images ranked highest to lowest from left to right. Green colors indicate correct and red for incorrect retrievals.